is a variable that has gaps between successive, possible values. Round proportions to three decimal places 6) A total of 360 American adults responded to a survey about their favorite movie genre. The table be summarizes their responses. low Horror Total 184 176 360 Romantic Action 53 Men Womer Total 98) 72 56101 8948 170 a) What proportion of the adults like to watch horror movie? (2 pts.) b) _What proportion of women like to watch romantic OR horror movie? (2 pts.) 565 c) What proportion of the adults like to watch action movie, given that they are men? (2 pts.)
Answer #1

a)

Out of 360 adults, 101 watch horror movie so

P(horror movie) = 101 /360 = 0.2806

Answer: 0.281

b)

From table:

P(romantic) = 89/360

Since these are mutually exclusive so

P(romantic or horror) =P(romantic)+P(horror) = 89/360 + 101/360 = 190/360= 0.5278

Answer: 0.528

c)

Out of 184 men, 98 watch action movie so

P(action | men) = 98 / 184 = 0.5326

Answer: 0.533